find four consecutive odd intergers such that twice the sum of the first two numbers is eight more than the sum of the last two numbers
Answer by checkley71(8405)

2(X+X+2)=(X+4)+(X+6)+8 OR 2X+2X+4=2X+10+8 OR 4X-2X=18-4 OR 2X=14 THEN X=7
SO THE NUMBERS ARE 7,9,11 & 13
PROOF 2(7+9)=11+13+8 OR 2*16=32 OR 32=32